# Repository for extended dark matter object constraints

### v1

@@ 1 Introduction
These objects, commonly called MACHOs, are normally created in the early universe, and their most popular example are primordial black holes (PBHs) Green and Kavanagh (2021); Carr et al. (2016). However, we will focus on those massive compact objects that do not have a black hole in their interior, so-called extended dark matter objects (EDOs). These are a popular dark matter candidate due to the different formation mechanisms they can have, which makes them very generic in principle, but has the downside of needing to be constrained case by case depending on their mass profile. Still, strong constraints have been placed using gravitational (micro-)lensing techniques Croon et al. (2020a, b), gravitational waves Croon et al. (2023), accretion of baryonic matter onto the CMB spectrum Bai et al. (2020); Croon and Sevillano Muñoz (2024) and dynamical heating of stars Graham and Ramani (2024).
@@ 2 Definition of EDO profiles
Another common assumption for both EDOs and PBHs alike is that for sub-fractions of DM ($f_{\rm DM}<1$), the density of objects in a galaxy still follows the DM halo profile. Relaxing this assumption could, in principle, either weaken or strengthen the constraints on DM sub-fractions.
@@ 3.1 User’s guide
This repository is an extension of the PBH constraints repository created by Bradley Kavanagh Kavanagh , so readers familiar with that repository will find most aspects to be similar. The main difference is that, contrary to PBHs, we need to specify the object’s mass profile and the set of radii we want to plot. As we can see, the first line allows the user to choose the mass profile using one of the different shapes defined in the previous section. Then, the next entry specifies the different displayed radii, where we need to input $r_{90}$ in powers of 10 Solar radii. The three numbers in the entry correspond to the lowest radius, highest radius and the difference between each plotted radii, respectively. Finally, the last part sets the constraints we want to add to the plot, where different columns change the bounds display as follows: Bound corresponds to the name as it appears in the files folder; [^2] Using the Lensing bound shows all lensing bounds, currently being EROS-2, OGLE-IV and Subaru-HSC, and gives the combination of these as a single constraint. linestyle allows the change of the plotting line style; x, y, and rotation set the position and inclination of the displayed text, defined by the last column. Contrary to PBHs bounds, we do not allow to change the colour of a bound since this degree of freedom will be used to plot multiple radii at once.
@@ 3.2 Contributing to the repository
Once a complete bounds folder is created, the instructions to add it to the repository can be found on github.com/SergioSevi/EDObounds. Having bounds for all different radii and aforementioned mass profiles is not necessary to contribute to this repository. As long as the bounds correspond to one of the mass profiles described earlier, they will automatically added to the plot to which they correspond. For example, in Figure 4, we can see that some bounds only exist for one of the mass profiles and others only for the smallest sizes. Additionally, if the set of constraints converges beyond a certain critical radius, the code will automatically apply these bounds to any smaller radii.
Figure 4: Example for generated plots with an incomplete set of bounds. On the left, we have an NFW sub-halo, where it is clear that the Lensing and CMB bounds do not cover the whole range of plotted radii due to the nature of constraints. On the left, in addition to this lack of constraints on Lensing and CMB, there are no bounds on dynamic heating as they have not yet been calculated for these objects.

### v2

@@ 1 Introduction
These objects, commonly called MACHOs, are normally created in the early universe, and their most popular example are primordial black holes (PBHs) Green and Kavanagh (2021); Carr et al. (2016). However, we will focus on those massive compact objects that do not have a black hole in their interior, which we call extended dark matter objects (EDOs). These are a popular dark matter candidate due to the different formation mechanisms they can have, which makes them very generic in principle, but has the downside of needing to be constrained case by case depending on their mass profile. Still, strong constraints have been placed using gravitational (micro-)lensing techniques Croon et al. (2020a, b), gravitational waves Croon et al. (2023), accretion of baryonic matter onto the CMB spectrum Bai et al. (2020); Croon and Sevillano Muñoz (2024) and dynamical heating of stars Graham and Ramani (2024). Depending on the underlying theory producing the EDOs, they can be constrained by their introduction of beyond the standard model physics, as is the case for axion stars Tkachev (2015), among others. However, here, we will only consider those constraints coming from their gravitational influence on our universe, which are guaranteed to be model-independent.
@@ 2 Definition of EDO profiles
Another common assumption for both EDOs and PBHs alike is that for sub-fractions of DM ($f_{\rm DM}<1$), the density of objects in a galaxy still follows the DM halo profile. Relaxing this assumption could, in principle, either weaken or strengthen the constraints on DM sub-fractions. Additionally, we will assume a monochromatic distribution of EDO masses for all cases. This approach may not represent the most realistic scenario in every context, but it avoids committing to any specific mass distribution, making it a useful middle ground for comparison. Moreover, as with PBHs, the bounds for other mass distribution functions can be derived from the monochromatic data Carr et al. (2017).
@@ 3.1 User’s guide
This repository is an extension of the PBH constraints repository created by Bradley Kavanagh Kavanagh , so readers familiar with that repository will find most aspects to be similar. The main difference is that, contrary to PBHs, we need to specify the object’s mass profile and the set of radii we want to plot. As we can see, the first line allows the user to choose the mass profile using one of the different shapes defined in the previous section. Then, the next entry specifies the different displayed radii, where we need to input $r_{90}$ in powers of 10 Solar radii. The three numbers in the entry correspond to the lowest radius, highest radius and the difference between each plotted radii, respectively. Finally, the last part sets the constraints we want to add to the plot, where different columns change the bounds display as follows: Bound corresponds to the name as it appears in the files folder; [^2] Using the Lensing bound shows all lensing bounds, currently being EROS-2, OGLE-IV and Subaru-HSC, and gives the combination of these as a single constraint. Similarly, using All will combine all bounds as a single constraint. linestyle allows the change of the plotting line style; x, y, and rotation set the position and inclination of the displayed text, defined by the last column. Contrary to PBHs bounds, we do not allow to change the colour of a bound since this degree of freedom will be used to plot multiple radii at once.
@@ 3.2 Contributing to the repository
Once a complete bounds folder is created, the instructions to add it to the repository can be found on github.com/SergioSevi/EDObounds. Having bounds for all different radii and aforementioned mass profiles is not necessary to contribute to this repository. As long as the bounds correspond to one of the mass profiles described earlier, they will automatically add to the plot to which they correspond [^3] In the case these bounds cannot be extended to the mass profiles described in Section 2, a new folder can be added to the repository for a different EDO shape, if necessary. However, we believe it is advantageous to generate all constraints for the same type of object, allowing for easier comparison across bounds.. For example, in Figure 4, we can see that some bounds only exist for one of the mass profiles and others only for the smallest sizes. Additionally, if the set of constraints converges beyond a certain critical radius, the code will automatically apply these bounds to any smaller radii.
Figure 4: Example for generated plots with an incomplete set of bounds. On the left, we have an NFW sub-halo, where it is clear that the Lensing and CMB bounds do not cover the whole range of plotted radii due to the nature of constraints. On the right, in addition to this lack of constraints on Lensing and CMB, there are no bounds on dynamic heating as they have not yet been calculated for these objects.